OpenStudy (anonymous):

Six less than the product of two consecutive positive integers is equal to six times their sum. Find the two integers please help

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

solve \\[n(n+1)-6=6(n+n+1)\]

OpenStudy (anonymous):

so it'll be n^2+n-6=6n+6n+6??/

OpenStudy (anonymous):

Correct. The first positive integer will be equal to n, and the second positive integer will be n + 1

OpenStudy (anonymous):

so do i have to factor it out, to get the answer for the first integer?